# Multi-Scale Cell Decomposition for Path Planning using Restrictive Routing Potential Fields

### v1

@@ Abstract
In burgeoning domains, like urban goods distribution, the advent of aerial cargo transportation necessitates the development of routing solutions that prioritize safety. This paper introduces Larp, a novel path planning framework that leverages the concept of restrictive potential fields to forge routes demonstrably safer than those derived from existing methods. The algorithm achieves it by segmenting a potential field into a hierarchy of cells, each with a designated restriction zone determined by obstacle proximity. While the primary impetus behind Larp is to enhance the safety of aerial pathways for cargo-carrying Unmanned Aerial Vehicles (UAVs), its utility extends to a wide array of path planning scenarios. Comparative analyses with both established and contemporary potential field-based methods reveal Larp’s proficiency in maintaining a safe distance from restrictions and its adeptness in circumventing local minima.
@@ I Introduction
As we look to the future, the role of aerial cargo delivery is expected to become central in the domain of urban goods transportation. Given its growing importance, the development of solutions that are both scalable and verifiably safe is crucial [1, 2, 3]. Safety is a paramount concern, and routing solutions are integral to this aspect. The flight paths chosen for unmanned aerial vehicles (UAVs) significantly influence the potential for aerial incidents. Consequently, a route’s function now extends beyond mere connectivity from origin to destination; it must also encompass considerations for safety.
This paper introduces a path planning framework tailored for restrictive routing within potential fields. It decomposes a potential field into multi-scale cells with an estimated maximum potential for their region and utilizes graph path planning algorithms to generate routes that overcome several of the issues with traditional artificial potential field methods. Initially conceived to bolster the safety and scalability of UAV-operated aerial cargo transport within urban air mobility (UAM), the algorithm’s utility extends beyond this scope. It is versatile enough to be implemented in any setting that aligns with the potential field standard to be discussed [4].
@@ I-A Background
A variety of distinct and diverse routing approaches have been proposed for the path planning of UAVs and similar vehicles in complex environments [5]. Two prominent categories are sampling-based and node-based approaches, which involve sampling or deconstructing the environment into networks of nodes or cells for navigation, respectively. Any sector containing an obstacle is excluded from consideration, and the remaining space is searched for optimal routing. Node-based and some sampling-based approaches often employ established graph path planning algorithms to determine routes after decomposing the feasible space. Notable algorithms in these categories include rapidly-exploring random trees (RTTs), artificial potential fields (APFs), A*, and their variants [6, 7, 8, 9, 10, 11, 12].
Another major category gaining popularity is mathematical model-based approaches. These methods formulate and apply mathematical expressions to determine optimal routes either through optimization or by solving numerical equations. A prominent approach in this category is mixed-integer linear programming (MILP) [13, 14].
Bio-inspired algorithms, as the name suggests, leverage intrinsic properties found in nature to design optimal routes. Ant colony optimization (ACO) and particle swarm optimization (PSO) have become standard algorithms within this category [15, 16, 17].
Reinforcement learning-based algorithms are a recent class of algorithms that have emerged from advancements in the field of machine learning. These methods fundamentally optimize a model for sequential decision-making in route search [18, 19, 20]. A subset of these methods has also been trained to work with potential fields for path planning [20]. However, due to the ‘black box‘ nature of data-driven models and their non-deterministic output, machine learning models are not considered for comparison in this research. Routes for cargo UAVs must be reproducible, scalable, and verifiable for practical application.
I-B Contributions
A significant portion of the works reviewed focus on grid-based environments, which may not accurately reflect real-world situations. Moreover, only a handful of these works actively incorporate variable proximity to restrictions as part of their route search. Our work stands out as one of the few to considers the safety of the generated routes and their potential to violate restrictions.
@@ I-B Contributions
Akin to our work, [12, 21] explore the application of potential fields for path planning, with [12] employing them specifically to quantify safety. These studies also utilize graph-based algorithms such as A* to navigate around local minima, a well-known issue associated with APFs methods. However, the methodology of [12] is confined to grid-based environments, and [21] primarily applies A* within a grid context to overcome local minima, which may lead to routes that are not optimized for distance.
Our work expands upon these concepts by considering both continuous and complex spaces within a unified potential field framework. We examine dynamic proximity to obstacles using multi-resolution cells and assign a safety guideline to routes in both complex and non-complex environments through the concept of potential for restriction violation.
@@ II-A Restrictive Routing Potential Fields
In [4], we explore the problem of last-mile UAVs air management and aerial cargo delivery in urban environments, and introduce a standard for restrictive routing based on a repulsive artificial potential field. Assuming a constant operating altitude, physical and virtual restrictions are designated as areas of high potential that diminishes the further an UAV (or agent) is from an obstacle; a behavior dictated by their repulsion matrix.
The standardization facilitates an uniform approach to potential field construction, enabling consistent replication and analysis across different routing algorithms. By leveraging the structured nature of the field, such algorithms can efficiently navigate through complex environments utilizing a common standard, avoiding obstacles, and minimizing the risk of restriction violations. The potential field delineated herein is henceforth denoted as a ‘restrictive routing potential field‘.
@@ II-B Standardized Potential Field Units
Under the proposed standard, a restrictive routing potential field is comprised of multiple building block units, each corresponding to a distinct type of restriction. Upon examining their definitions, a set of intrinsic properties emerges: a repulsion vector $\bar{x}$, a scaled squared distance $\tilde{d}\,^{2}(x)$, and a potential evaluation $\sigma(x)$. The repulsion vector $\bar{x}$ indicates the direction and magnitude of the force repelling from restricted areas. The scaled squared distance $\tilde{d}\,^{2}(x)$ reflects the scaled proximity to a point $x$, modulated by the unit’s repulsion matrix $A$. Finally, the potential $\sigma(x)$ denotes the field’s influence at point $x$, effectively reflecting the potential for restriction violation.
The repulsion vectors for the fundamental units and collection of them are detailed in Table I.
@@ III Methodology
Given the properties of the standardized potential field, we introduce Larp (Last-mile restrictive path planning) as a routing framework for safe path planning. It consists of three core stages: first, an primary algorithm that decomposes a field into distinct multi-scale cells; secondly, the formation of a routing network using the cells and an adjacency detection algorithm; lastly, it traverses the graph to determine a suitable route. The stages can be observed in Fig. 1. An official implementation of Larp can also be found on our Github repository at https://github.com/wzjoriv/Larp.
@@ III-A Multi-Scale Cell Decomposition
Drawing inspiration from adaptive quad tree cell decomposition, the first phase consists of partitioning a potential field into a multitude of cells, each varying in size. These cells are then correlated with their respective potential for infringing upon nearby restrictions. The stratagem for cell subdivision is calibrated, ensuring that cells proximal to obstacles are proportionately smaller, thereby augmenting routing precision and safety. Fig. 1(b) illustrates this decomposition for a walled room scene. The methodology behind cell decomposition is expounded in Algorithms 1 and 2.
@@
minimum cell size $n_{min}$; maximum cell size $n_{max}$
Cells containing restriction have highest potential
function Build($x$, $n$, $U$)
     $quad\leftarrow QuadNode(x,n)$
     $zones\leftarrow$ map with default farthest zone
     if $|U|>0$ then
         $zones\leftarrow$ ApproxDistanceZones($x$, $n$, $U$)
     end if
     $quad.zone\leftarrow\min(zones)$triangleright Farthest zone if empty
     if $n\leq n_{max}$ then
         if $n\leq n_{min}$ or $quad.zone$ is farthest zone then
              Mark $quad$ as leaf
              return $quad$
         else if $quad.zone>0$ then
@@
              $\bar{C}\leftarrow\{c=x+\frac{n}{\sqrt{2}\,|\bar{x}_{u}(x)|}\,\bar{x}_{u}(x)\,|\,u\in\,\bar{U}\}$
              $bd_{l}\leftarrow boundary\_egde[quad.zone]$
              for each point $c$ in $\bar{C}$ do
                  if given any $u\in\bar{U}$, $\tilde{d}_{u}^{2}(x)<bd_{l}$ then
                       Mark $quad$ as leaf
                       return $quad$
                  end if
              end for
         end if
     end if
[…]
     for each quad $q$ of a cell do
         $quad.child[q]\leftarrow$ Build($q.center$, $n/2$, $U$)
     end for
     return $quad$
end function
$root\leftarrow$ BUILD(field.center, field.size, all units in field)
Algorithm 2 Larp – Estimate Boundary of Cell
Sorted list of boundaries $bds$ in ascending order
function ApproxDistanceZones($x$, $n$, $U$)
     $zones\leftarrow$ map with default further zone
     for each unit $u$ in $U$ do
         if $d_{u}^{2}(x)\leq\frac{n^{2}}{2}$ then
              $zones[u]\leftarrow 0$
         end if
     end for
     for each unit $u$ in $U$ but not already in $zones$ do
         $dist=\tilde{d}_{u}^{2}(x)$
         Assign $zones[u]$ into bin given $bds$ and $dist$
     end for
     return $zones$
end function
@@ III-B Routing network
To streamline path planning within the designated space, a network graph is constructed subsequent to the potential field’s decomposition. The process employs a divide-and-conquer approach to traverse the quad tree, facilitating the identification of neighboring nodes for any given leaf quad node. The intricacies of the field are distilled into a simplified network, where each graph node correlate with a leaf cell and its upper bound potential for the vicinity. Fig. 1(c) provides a visual representation of this network graph.
@@ III-C Path Planning
Upon the establishment of the network, the implementation of any graph routing algorithm that simultaneously minimizes route distance and accumulated potential is feasible. For this study, we employ a modified variant of the renowned A* algorithm, wherein the distance function is redefined as:
@@
where $q_{a}$ and $q_{b}$ denote the quadrant being traversed from and to, respectively; $s(q)$ represents a hyper-parameter scale transformation function that modulates the safety distance; and $q_{a}.\text{center}$ signifies the centroid of cell $q_{a}$. A route generated for the walled room scene is depicted in Fig. 1(d). Subsequent to determining a routing solution, line segments are integrated to connect the initial location and the destination with the discovered route within the network.
@@ III-D Safety Validation
Prior to assessing the safety of a route, it is imperative to define what constitutes a route within a potential field.
@@
The extended duration of a route implies a prolonged presence within the field, potentially elevating the risk of violation occurrences.
@@ IV-A3 Modified Artificial Potential Field (M-APF)
IV-B Results
In evaluating these algorithms, we explored three distinct scenarios of path planning.
@@ IV-B1 Scenario 1 (Unobstructed Path)
Algorithm Goal Found Route Distance Route Area Average Potential Highest Potential
@@ IV-B2 Scenario 2 (Obstructed Path and Near-Obstacle Goal)
Algorithm Goal Found Route Distance Route Area Average Potential Highest Potential
@@ IV-B3 Scenario 3 (Walled-Off Path)
. Algorithm Goal Found Route Distance Route Area Average Potential Highest Potential PM 33.32 18.6015 0.5583 0.9529 APF 42.83 26.3972 0.6163 0.9075 APF (*) 80.37 53.9457 0.6712 0.8075 M-APF 38.57 1.8687 0.0485 0.0674 Larp (Our) 163.8382 14.8769 0.0908 0.3453

### v3

@@ Abstract
In burgeoning domains such as urban goods distribution, the advent of aerial transportation necessitates the development of routing solutions that prioritize safe navigation. This paper introduces Larp, a novel path planning and navigation framework that leverages the concept of repulsive potential fields as continuous cost maps to forge safe routes. The algorithm achieves it by segmenting the potential field into a hierarchy of cells, each with a designated risk zone determined by the proximity of obstacles. The meshing allows the airspace to be partitioned based on an area’s potential for restriction violations, enabling navigation that is aware of these risks. While the primary impetus behind Larp is to enhance the safety of aerial pathways for Unmanned Aerial Vehicles (UAVs) in urban air mobility, its utility extends to a wide array of routing scenarios. Comparative analyses with both established and contemporary potential field-based methods reveal Larp’s proficiency in maintaining a safe distance from restrictions and its adeptness in circumventing local minima. Additionally, large-scale aerial path planning of Austin, TX demonstrates Larp’s capability to be implemented at a large scale.
Keywords:
potential field; cost maps; air traffic management; path planning; unmanned aerial vehicle; urban air mobility
@@ I Introduction
As we look to the future, the role of aerial cargo delivery is expected to become central in the domain of urban goods transportation. Given its growing importance, the development of solutions that are both scalable and verifiably safe is crucial [1, 2, 3, 4]. Routing solutions are integral to this aspect. The flight paths chosen for unmanned aerial vehicles (UAVs) significantly influence the potential for aerial incidents. Consequently, a route’s function now extends beyond mere connectivity from origin to destination; it must also encompass considerations for safety.
This paper introduces Larp, a path planning and navigation framework tailored for restrictive routing within potential fields as cost maps. The framework decomposes a potential field into multi-scale cells, each with an estimated maximum potential for its region, and utilizes graph path planning algorithms to generate routes that overcome several issues associated with traditional artificial potential field methods. Although primarily conceived to bolster the safety and scalability of UAV-operated aerial cargo transport within urban air mobility (UAM), the algorithm’s utility extends beyond this scope. It is versatile enough to be implemented in any setting that aligns with the potential field standard to be discussed [5].
@@ I-A Background
Cost maps in path planning describe a routing scheme in which sectors of a map are assigned a cost or penalty for being in a particular region, and routes are formulated to minimize the accumulated cost of navigation [6, 7]. These maps are utilized in mobile robotics for path planning with risk assessment to effectively navigate obstacles [7, 8]. Urban air mobility, a multifaceted endeavor, necessitates the consideration of various factors, including obstacle clearance, noise pollution, privacy concerns, and ground risks [4, 9]. Cost maps serve as a powerful tool to represent general restriction violations by abstracting the complexity of risk assessment into a spatial cost metric. They have shown significant promise for UAV path planning based on risk assessment [6, 10, 11, 12]. Demonstrated benefits include reducing risks in metropolitan areas, integrating multiple risk factors into a single cost map, and aiding in prioritizing search areas during search-and-rescue missions. Another critical avenue explored is the inclusion of airspace occupancy in cost maps, facilitating routing that minimizes conflicts with other aerial vehicles operating in the same airspace [13, 14].
Artificial potential fields define a continuous mathematical system with attractive and repulsive forces that move particles towards a goal. Areas to avoid are classified as having high repulsive potential that decays with distance from the particle. Traditionally, path planning with potential fields involves simulating particles influenced by the repulsive field while being attracted towards the goal via an attractive force [15, 16, 17, 18, 19, 20, 21, 22, 23]. These fields are utilized for a variety of autonomous vehicles. In dynamic environments, potential fields have been used to guide the navigation of autonomous underwater vehicles (AUVs) with moving obstacles [24]. Potential fields have also shown promise for UAV path planning in both simulation and real-world environments [25, 26, 15, 16]. Their use has led to increased safety by ensuring a maintained distance from obstacles and, at times, faster path searches due to the system’s simplicity and its heuristic nature. However, research on artificial potential fields for UAVs is limited [27]. The main drawback of potential field methods is that they may form local minima that impede agents from reaching their destination. From one perspective, the repulsive component of a potential field can be classified as a class of continuous cost maps, defined in a space where any point in the field has a potential that must be minimized. We seek to combine this concept with variable resolution cost maps to generate safer routes in complex continuous spaces.
@@
Similar to our work, [15, 16] investigate the application of potential fields for path planning, with [15] specifically utilizing them to quantify safety. These studies also employ graph-based algorithms such as A* to navigate around local minima, a well-known issue associated with artificial potential field (APF) methods. However, the methodology of [15] is confined to fixed grid-based environments, and [16] primarily applies A* within a grid context to overcome local minima, which may lead to routes that are not optimized for distance. While cell decomposition has widespread use for environment representation in UAV operations, few studies consider artificial potential fields [27]. We propose a framework for cell decomposition that relies on artificial potential fields as cost maps. This approach allows for the examination of variable proximity to obstacles using an analytical method for multi-resolution cell decomposition and the assignment of safety guidelines to routes based on the expected accumulated potential. Moreover, a significant portion of the reviewed works focus on grid-based environments, which may not accurately reflect real-world situations. Our work operates in both continuous and complex spaces through a standardized format.
@@ II-A Restrictive Routing Potential Fields
In [5], we explore the problem of last-mile UAVs air traffic management and aerial cargo delivery in urban environments, and introduce a standard for restrictive routing using cost map based on repulsive artificial potential fields. Assuming a constant operating altitude, physical and virtual restrictions are designated as areas of high potential that diminishes the further an UAV (or agent) is from an obstacle; a behavior dictated by their repulsion matrix.
The standardization facilitates an uniform approach to cost map construction, enabling consistent replication and analysis across different routing algorithms. By leveraging the structured nature of the potential field, such algorithms can efficiently navigate through complex environments utilizing a common standard, avoiding obstacles, and minimizing the risk of restriction violations. Additionally, it serves as a foundation for urban air traffic management and to analyze and study bottlenecks in the feasible airspace. The potential field delineated herein is henceforth denoted as a ‘restrictive routing potential field‘.
@@ II-B Standardized Potential Field Units
Under the proposed standard, a restrictive routing potential field is comprised of multiple building block units, each corresponding to a distinct type of restriction. Upon examining their definitions, a set of intrinsic properties emerges: a repulsion vector $\bar{x}$, a scaled squared distance $\tilde{d}\,^{2}(x)$, and a potential evaluation $\sigma(x)$. The repulsion vector $\bar{x}$ encodes the direction from and proximity to the restricted areas. The scaled squared distance $\tilde{d}\,^{2}(x)$ reflects the weighted proximity to a point $x$, modulated by the unit’s repulsion matrix $A$. Finally, the potential $\sigma(x)$ denotes the field’s influence at point $x$, effectively reflecting the potential for restriction violation.
The repulsion vectors for the fundamental units and collection of them are detailed in Table I. Fig. 1 illustrate the repulsion vectors for an arbitrary set of obstacles with respect to an arbitrary point. The magnitude of each vector indicates proximity to the obstacle and the direction provides a guide towards moving away from it.
@@
Figure 1: Illustration of the repulsion vectors with respect to a location.
@@ III Methodology
Given the properties of the standardized potential field, we introduce Larp (Last-mile restrictive path planning) as a routing framework for safe path planning and navigation. It consists of three main stages: first, a core algorithm that decomposes a field into distinct multi-scale cells; secondly, the formation of a routing network using the cells and their adjacency; lastly, the search for a feasible, suitable, and safe route. The stages can be observed in Fig. 2.
@@
Note that an official implementation of Larp can also be found on our Github repository at https://github.com/wzjoriv/Larp.
@@ III-A Multi-Scale Cell Decomposition
Drawing inspiration from adaptive quad tree cell decomposition, the first and main phase consists of partitioning a potential field into a multitude of cells, each varying in size. These cells are then correlated with their respective potential for infringing nearby restrictions. The stratagem for cell subdivision is calibrated, ensuring that cells proximal to obstacles are proportionately smaller, thereby augmenting routing precision and safety. Fig. 2(b) illustrates this decomposition for a walled room scene. The methodology behind cell decomposition is expounded in Algorithms 1 and 2.
III-A1 Cells Quad Tree Structure
Algorithm 1 consists of a recursively subdividing a field into quadrants (i.e., quad or cell) and assigning a restriction zone based on the area potential for violating a restriction. This forms a tree structure with quad cells as node. The tree allows for quick search of any area and the restrictions in it. An example of the quad tree structure in presented in Fig. 3 for a simple field. Initially, at the start of the algorithm, an empty node is initialize for the current observed quadrant. Then, via Algorithm 2, and the highest (i.e., closest) risk zone is assigned from a list of maximum potential limits. Once a zone has been assigned, it is determined wether the subdivision should stop or continues. It stops if we have assigned the further zone allowed or the minimum desired cell size has been reached. After that, we remove any obstacle from consideration that is to far from the quadrant and subside again for each quadrant inside the current sector. The leaves of the generated tree present the multi-scale cell decomposition of the potential field and the individual restriction zone.
Figure 3: Quad tree for a simple potential field.
@@ III-A1 Cells Quad Tree Structure
1:function Build($x$, $n$, $U$)
2:     $quad\leftarrow QuadNode(x,n)$
3:     $zones\leftarrow$ map with default farthest zone
4:     if $|U|>0$ then
5:         $zones\leftarrow$ ApproxObstaclesZones($x$, $n$, $U$)
6:     end if
7:     $quad.zone\leftarrow\min(zones)$triangleright Farthest zone if empty
8:     if $n\leq n_{max}$ then
9:         if $n\leq n_{min}$ or $quad.zone$ is farthest zone then
10:              Mark $quad$ as leaf
11:              return $quad$
12:         end if
13:     end if
@@
15:     for each quad $q$ of a cell do
16:         $quad.child[q]\leftarrow$ Build($q.center$, $n/2$, $U$)
17:     end for
18:     return $quad$
19:end function
20:$root\leftarrow$ BUILD(field.center, field.size, all units in field)
III-A2 Repulsion Vectors and Potential of an Area
The heart of the cell decomposition is estimating the maximum potential of any cell in the potential field and assigning a restriction zone. This is done by Algorithm 2. By relying on the intrinsic properties of the standardized potential field, an maximum potential to violate restrictions can analytically be determined. Initially, we calculate a proxy of the square distance of each obstacle unit with respect to the center of a quad. If the distance is less than the radius of the circumscribed circle of the quad, then the obstacle is assigned zone 0 (i.e., an obstacle exist in the area). Trivially, this ensures that if an obstacle is in the area, it is classified as being in zone 0. For the other zones, we must rely on the scaled distance to the obstacles. First, we use the repulsion vector of each obstacle to estimate a location $c$ of where the obstacle is closest to the area. Then, we sample the potential at this location for each obstacle individually and assign it to a zone based on how close they are to the value $1.0$. Each zone assignment is based the potential of any obstacle to be violated by an agent in the area (or indirectly their weighted proximity of the sector to the restriction).
Algorithm 2 Larp – Estimate Maximum Potential Zone of a Quad
1:Sorted list of boundaries $bds$ in descending order
2:function ApproxObstaclesZones($x$, $n$, $U$)
3:     $zones\leftarrow$ map with default farthest zone
4:     for each unit $u$ in $U$ do
5:         if $d_{u}^{2}(x)\leq\frac{n^{2}}{2}$ then
6:              $zones[u]\leftarrow 0$
7:         end if
8:     end for
9:     for each unit $u$ in $U$ but not already in $zones$ do
10:         $c=x-\frac{n}{\sqrt{2}\,\|\bar{x}_{u}(x)\|}\,\bar{x}_{u}(x)$
11:         $p=\exp{\left(\tilde{d}_{u}^{2}(c)\right)}$
12:         Assign $zones[u]$ into bin given $bds$ and $p$
13:     end for
14:     return $zones$
15:end function
Fig. 4 illustrates how an area maximum potential (or risk) zone can be approximated. First, the repulsion vector of each obstacle is flipped and clamped to the magnitude of the diagonal distance from the quad’s center to to its edge. In the figure, they are drawn from the center of the area and colored based on the maximum potential assigned. The maximum potential ranges for this example are ticked in the color bar. As the potential for the obstacle to the right of the center is the closest to $1.0$, the entire circular area is estimated to have an maximum potential of $0.8$. As a subset of the circular area, the quad is also estimated to have an maximum potential upper limit of $0.8$.
Figure 4: Estimating the maximum potential of an given area in the potential field.
@@ III-B Routing Network
To streamline path planning within the designated space, a network graph is constructed subsequent to the potential field’s decomposition. The process employs a divide-and-conquer approach to traverse the quad tree, facilitating the identification of neighboring nodes for any given leaf quad node. The intricacies of the field are distilled into a simplified network, where each graph node correlate with a leaf cell and its maximum potential for the vicinity. Fig. 2(c) provides a visual representation of this network graph.
@@ III-C Path Planning
Upon the establishment of the network, the implementation of any graph routing algorithm that simultaneously minimizes route distance and accumulated potential is applicable. For this study, we employ a modified variant of the renowned A* algorithm, wherein the distance function is redefined as:
@@
where $q_{a}$ and $q_{b}$ denote the quadrant being traversed from and to, respectively; $s(q)$ represents scale transformation function that modulates the penalty of going to a new sector given its maximum potential; and $q_{a}.\text{center}$ signifies the centroid of a cell $q_{a}$. A route generated for the walled room scene is depicted in Fig. 2(d). Subsequent to determining a routing solution, line segments are integrated to connect the initial location and the destination with the discovered route within the network.
@@ III-D Safety Validation
It is imperative to establish a commonality for defining what constitutes a route within a potential field to assess effectiveness of the routing solutions.
@@
The extended duration of a route implies a prolonged presence within the field, elevating the risk of violation occurrences.
@@ IV-A3 Modified Artificial Potential Field (M-APF)
IV-B Results: Routes in a small room
In evaluating these algorithms, we explored three distinct scenarios of path planning in a small room.
@@ IV-B1 Scenario 1 (Unobstructed Path)
Algorithm Goal Found Distance ($R_{d}$) Potential ($R_{A}$) $R_{avg}$ Highest Potential
@@ IV-B2 Scenario 2 (Obstructed Path and Near-Obstacle Goal)
Algorithm Goal Found Distance ($R_{d}$) Potential ($R_{A}$) $R_{avg}$ Highest Potential
@@ IV-B3 Scenario 3 (Walled-Off Path)
. Algorithm Goal Found Distance ($R_{d}$) Potential ($R_{A}$) $R_{avg}$ Highest Potential PM 33.32 18.6015 0.5583 0.9529 APF 42.83 26.3972 0.6163 0.9075 APF (*) 80.37 53.9457 0.6712 0.8075 M-APF 38.57 1.8687 0.0485 0.0674 Larp (Our) 163.8382 14.8769 0.0908 0.3453
@@
IV-C Results: UAV routes in Austin, TX
Larp is capable of generating safe city-level routes. To test its capability, UAV routes for Austin, TX where generated using OpenStreetMap buildings data. it is assumed that all buildings are tall enough to interfere with UAVs travel, while anything outside of the area does not. Fig. 8 and 9 explore two similar but distinct scenarios involving urban air mobility.
In Fig. 8, travel within an urban environment is explored, specifically focusing on safe route generation inside a dense neighborhood with high buildings. As shown, Larp has the ability to generate routes that maintain a safe distance from building boundaries (colored deep red with a potential of one) wherever possible. It only enters a building’s airspace when required to perform the delivery. Additionally, routes can adapt to crevices created by a number of buildings. Note that the scale of the axes is in meters, and the policy for proximity to buildings can be adjusted during path planning by the penalty function. For this demo, we have chosen to set an upper potential for restriction violation at 0.2, only getting closer when necessary without violating boundary restrictions.
(a) Cell decomposition of dense neighborhood.
(b) Initial delivery points.
(c) Routes generated by Larp.
Figure 8: Routes among high buildings in downtown Austin, TX.
Fig. 9 presents the results when the algorithm is expanded to a larger scale, including rural or unobstructed areas. In this scenario, the surrounding areas of Austin transition from a high dense urban sector to an unobstructed space. As observed, Larp generates coarse paths when away from buildings and granular paths when near them. The coarse paths allow for easier straight flights, which are convenient for long-distance travel without drastic maneuvering. When near a building, the granular, smoother, and more precise routes allow for safer navigation around the buildings. For the demonstration, and as illustrated, the routes for UAVs avoid buildings wherever possible. However, to ensure deliveries to buildings, we have opted to allow entrance into a building’s airspace if necessary for delivery, as also seen in the previous scenario.
Figure 9: Routes surrounding downtown Austin, TX.
The source code to generate the routes near Austin, TX or any city can be found on our Github repository at https://github.com/wzjoriv/Larp. A GeoJSON file of the city’s buildings is needed to run it. It can be obtained via OpenStreetMap or similar map providers.